# Feng-Yu Wanga), Chenggui Yuanb), Xiao-Yu Zhaoa)

### v1

@@ 2. $(2)$
Combining this with (3.27) and applying Holder’s inequality, we find constants $c=c(p,q,\varepsilon)>0$ and $\theta=\theta(p,q)>0$ such that
$\mathbb{E}\left[\exp\left(\int_{0}^{t\land\tau_{n}}|f_{s}|(X(s))\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]$
$=$ $\mathbb{E}_{\mathbb{Q}_{n}}\left[R_{n}(T)^{-1}\exp\left(\int_{0}^{t\land\tau_{n}}|f_{s}|(X(s))\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]$
$\leq$ $\left(\mathbb{E}\big{[}R_{n}(T)^{-(\tilde{p}-1)}\big{|}\mathscr{F}_{0}\big{]}\right)^{\frac{1}{\tilde{p}}}\left(\mathbb{E}_{\mathbb{Q}_{n}}\left[\exp\left(\frac{\tilde{p}}{\tilde{p}-1}\int_{0}^{t\land\tau_{n}}|f_{r}|(X(s))\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]\right)^{\frac{\tilde{p}-1}{\tilde{p}}}$
@@
By Lemma 3.5 and (3.21), we have the following corollary which is a key in the proofs of Theorem 3.1 and Theorem 3.2.
@@ 1. $(1)$
    $\begin{split}&\mathbb{E}\left[\exp\left(\int_{s}^{t}|f_{u}(X(u))|\|X_{u}\|_{\tau}^{\alpha}\text{\rm{d}}u\right)\bigg{|}\mathscr{F}_{s}\right]\leq\exp\left[\varepsilon{\|X_{s}\|}_{\tau}^{2\alpha}+k\left(1+\|f\|^{\theta}_{\tilde{L}^{p}_{q}(s,t)}\right)\right],\\ &\qquad\ 0\leq s\leq t\leq T,\ f\in\tilde{L}^{p}_{q}(T).\end{split}$ (3.29)
@@ 2. $(2)$
By (3.2), for any $n\geq 1$, $b^{(1)}_{n}$ is bounded and satisfies $(A_{2}^{\prime})$. So, Lemma 3.3 ensures that the SDE
@@
For $\lambda\geq\lambda_{0}$ and $u$ solving the PDE (3.12) such that (3.14) holds, let
@@
By $(A_{1})$, $(A_{2})$, (3.14) and Itô’s formula, we find a constant $c_{1}>0$ such that
@@
Let $\Theta(t,x):=x+u(t,x)$ for $u$ solving the PDE (3.12) such that (3.14) holds, and denote
@@
By $(A_{1})$, $(A_{2}^{\prime})$, (3.14) and Lemma 3.4, we find a constant $k_{0}>0$ such that
@@
3.2 Asymptotic log-Harnack inequality and applications
Note that when the SDE (3.1) is time-homogenous and monotone, the asymptotic log-Harnack inequality has been derived in [5, Theorem 3.1] for the Markov semigroup
  1. $(B)$
    where ${\bf 0}\in\mathscr{C}_{\tau}$ with ${\bf 0}(r)=0,r\leq 0.$
Theorem 3.8.
[…]
$P_{t}\log f(\eta)\leq\log P_{t}f(\xi)+c\|\xi-\eta\|_{\tau}^{2}+c\text{\rm{e}}^{-\tau_{0}t}\|\nabla\log f\|_{\infty}\|\xi-\eta\|_{\tau}.$
Corollary 3.9.
In the situation of Theorem 3.8, the following assertions hold.
  1. $(1)$
  2. $(2)$
  3. $(3)$
    $\delta(\xi,A):=\liminf_{t\rightarrow\infty}P_{t}(\xi,A)>0.$
@@ 3. $(3)$
According to Theorem 3.1 and Corollary 3.7, under the assumptions $(H_{1})$$(H_{3})$, for any $\mu\in\mathcal{C}_{k}^{\gamma}$, the SDE
@@ 4.1 Proof of (a)
It suffices to prove (4.3) for $k>0$ and large $N$. By Lemma 3.6, $(H_{1})$-$(H_{3})$ implies that for some constant $c_{1}>0$ we have
@@
$\sup_{\mu\in\mathcal{C}_{k}^{\gamma}}\mathbb{E}\bigg{(}\int_{0}^{T}|f_{0}(s,X^{\mu}(s))|\text{\rm{d}}s\bigg{)}^{k}<\infty.$
Combining this with $(H_{1})$-$(H_{3})$, $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau})$, the Itô formula and the BDG inequality, we find constants $c_{2},c_{3}>0$ such that
@@
Using Grönwall’s lemma and noting that $\mu\in\mathcal{C}^{\gamma,N}_{k}$, we find a constant $c_{5}$ such that
$\begin{split}&\mathbb{E}\left[1+\|X_{t}\|_{\tau}^{k}\right]\leq c_{5}+c_{5}\int_{0}^{t}\text{\rm{e}}^{-k\tau(t-s)}\|\mu_{s}\|_{k}^{k}\text{\rm{d}}s\\ &\leq c_{5}+c_{5}N\text{\rm{e}}^{Nt}\int_{0}^{t}\text{\rm{e}}^{-N(t-s)}\text{\rm{d}}s\leq 2c_{5}\text{\rm{e}}^{Nt},\ \ t\in[0,T].\end{split}$ (4.6)
@@
$\begin{split}&\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\left(1+\|\Phi_{t}^{\gamma}\mu\|_{k}^{k}\right)=\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\mathbb{E}\left(1+\|X_{t}\|_{\tau}^{k}\right){\leq N},\quad N\geq N_{0},\ \mu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$
@@
$\begin{split}&U_{t}:=\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}\|X_{s}\|_{\tau}^{k}\right)\right]\leq c_{6}+c_{6}\mathbb{E}\left(\int_{0}^{t}\left\{\text{\rm{e}}^{2\tau s}\|X_{s}^{\mu}\|_{\tau}^{2}+\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\right\}\text{\rm{d}}s\right)^{\frac{k}{2}},\\ &\leq c_{7}+c_{7}\left(\int_{0}^{t}\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}+c_{7}\mathbb{E}\left\{\left[\sup_{s\in[0,t]}\text{\rm{e}}^{k\tau s}\|X_{s}\|_{\tau}^{k}\right]^{1-\frac{k}{2}}\left(\int_{0}^{t}\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\text{\rm{d}}s\right)^{\frac{k}{2}}\right\}\\ &\leq c_{7}+c_{7}\left(\int_{0}^{t}\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}+\frac{1}{2}U_{t}+c_{8}\int_{0}^{t}U_{s}\text{\rm{d}}s,\ \ t\in[0,T].\end{split}$
By Grönwall’s lemma, there exist constants $c_{9},c_{10}>0$ such that for any $\mu\in\mathcal{C}^{\gamma,N}_{k}$,
$\begin{split}&\mathbb{E}\left(1+\|X_{t}\|_{\tau}^{k}\right)\leq c_{9}+c_{9}\left(\int_{0}^{t}\text{\rm{e}}^{-2\tau(t-s)}\|\mu_{s}\|_{k}^{2}\text{\rm{d}}s\right)^{\frac{k}{2}}\\ &\leq c_{9}+c_{9}Ne^{Nt}\left(\int_{0}^{t}\text{\rm{e}}^{-2N(t-s)}\text{\rm{d}}s\right)^{\frac{k}{2}}\leq c_{9}+c_{10}N^{1-\frac{k}{2}}\text{\rm{e}}^{Nt},\ \ t\in[0,T].\end{split}$ (4.7)
@@
$\begin{split}&\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\left(1+\|\Phi_{t}^{\gamma}\mu\|_{k}^{k}\right)=\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\mathbb{E}\left(1+\|X_{t}\|_{\tau}^{k}\right)\\ &\leq c_{9}+c_{10}N^{1-\frac{k}{2}}\leq N,\ \ \mu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$
@@ 4.2 Proof of (b)
By Theorem 3.1, $(H_{1})$-$(H_{3})$, (4.3) and $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{k}(\mathscr{C}_{\tau})$, we find a constant $C_{0}(N)>0$ such that
@@
Next, by $(H_{1})$ and $(H_{2})$, we find a constant $C_{1}(N)>0$ such that for any $\mu,\nu\in\mathcal{C}^{\gamma,N}_{k}$,
@@
By $(H_{2})$ and Girsanov’s theorem, for any $t\in[0,T]$,
@@
$\begin{split}&\mathbb{E}\left(|R_{t}-1|^{2}\big{|}\mathscr{F}_{0}\right)\leq\text{\rm{e}}^{C_{1}(N)\int_{0}^{t}H(s)\left(1\wedge\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\right)\text{\rm{d}}s}\\ &\leq c_{2}(N)\left(\int_{0}^{t}H(s)\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\text{\rm{d}}s\right)^{\frac{1}{2}}\\ &\leq C_{3}(N)\text{\rm{e}}^{\theta t}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu)\left(\int_{0}^{t}H(s)\text{\rm{e}}^{-2\theta(t-s)}\text{\rm{d}}s\right)^{\frac{1}{2}},\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$ (4.11)
@@
where $\mathscr{L}_{X_{t}^{\mu}|\mathbb{Q}_{t}}$ stands for the distribution of $X_{t}^{\mu}$ under $\mathbb{Q}_{t}$. Thus, by (4.8), (4.9), (LABEL:N1) and Hölder’s inequality, we find constants $C_{4}(N)>0$ such that
@@
Noting that
@@
Let $k>1$. Let $\Theta(t,\cdot):=id+u(t,\cdot)$ for $u$ solving (3.12) such that (3.14) holds. Let
@@
$\text{\rm{d}}X^{\nu}(t)=b(t,X_{t}^{\nu},\nu_{t})\text{\rm{d}}t+\sigma(t,X^{\nu}(t))\text{\rm{d}}W(t),\ \ t\in[0,T]$ (5.3)
@@
Combining this with (3.17), Itô’s isometry and Yong’s inequality, we find a constant $c_{1}>0$ such that
@@
$\bar{\zeta}_{s}:=\left\{\sigma^{*}(\sigma\sigma^{*})^{-1}\right\}(s,X^{\mu}(s))\left[b(s,X^{\mu}_{s},\mu_{s})-b(s,X^{\mu}_{s},\nu_{s})\right],$
@@
$|\bar{\zeta}_{s}|\leq c_{1}\mathbb{W}_{2}(\mu_{s},\nu_{s})\leq c_{2}\text{\rm{e}}^{c_{2}s}\mathbb{W}_{2}(\mu,\nu)$ (5.6)
@@
$\text{\rm{d}}X^{\mu}(s)=b(s,X^{\mu}_{s},\nu_{s})\text{\rm{d}}s+\sigma(s,X^{\mu}(s))\text{\rm{d}}\bar{W}(s),\ \ \mathscr{L}_{X^{\mu}_{0}}=\mu,\ s\in[0,t].$ (5.7)
@@
$\text{\rm{d}}Y(s)=$$\{b(s,Y_{s},\nu_{s})+\kappa\sigma(s,Y(s))\sigma(s,X^{\mu}(s))^{-1}(X^{\mu}(s)-Y(s))\}\text{\rm{d}}s$ (5.8)
$+\sigma(s,Y(s))\text{\rm{d}}\bar{W}(s),\ \ s\in[0,t],\ \ Y_{0}=X_{0}^{\nu}.$
@@
$\tilde{\zeta}_{s}:=\kappa\sigma(s,X^{\mu}(s))^{-1}(X^{\mu}(s)-Y(s)),$
@@
$\text{\rm{d}}Y(s)=b(s,Y_{s},\nu_{s})\text{\rm{d}}s+\sigma(s,Y(s))\text{\rm{d}}\tilde{W}(s),\ \ s\in[0,t],\ \ Y_{0}=X_{0}^{\nu},$ (5.9)
which together with the uniqueness of (5.3) derive that $\mathscr{L}_{Y_{t}|\mathbb{Q}}=\mathscr{L}_{X_{t}^{\nu}}$. Moreover, if we choose $\mathscr{F}_{0}$-measurable random variables $X_{0}^{\mu}$ and $X_{0}^{\nu}$ on $\mathscr{C}_{\tau}$ such that
@@
[5, Lemma 3.3] and $(H^{\prime})$ imply that for any $p>0$, and $\tau_{0}\in(0,\tau)$, it holds
@@
for some constant $c>0$. Therefore, applying Young’s inequality in [2, Lemma 2.4],
@@
$\leq c_{3}\text{\rm{e}}^{c_{3}t}\mathbb{W}_{2}(\mu,\nu)^{2}+c_{3}\int_{0}^{t}\mathbb{E}_{\mathbb{Q}}\|X^{\mu}_{s}-Y_{s}\|_{\tau}^{2}\text{\rm{d}}s\leq c_{4}\text{\rm{e}}^{c_{4}t}\mathbb{W}_{2}(\mu,\nu)^{2}.$

### v2

@@ 2. $(2)$
Combining this with (4.22) and applying Hölder’s inequality, we find constants $c=c(p,q,\varepsilon)>0$ and $\theta=\theta(p,q)>0$ such that
$\mathbb{E}\left[\exp\left(\int_{0}^{t}\left|f^{n}_{s}(X(s\wedge\tau_{n}))\right|\,\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]$
$=$ $\mathbb{E}_{\mathbb{Q}_{n}}\left[R_{n}(T)^{-1}\exp\left(\int_{0}^{t}|f^{n}_{s}(X(s\wedge\tau_{n}))|\,\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]$
$\leq$ $\left(\mathbb{E}\big{[}R_{n}(T)^{-(\tilde{p}-1)}\big{|}\mathscr{F}_{0}\big{]}\right)^{\frac{1}{\tilde{p}}}\left(\mathbb{E}_{\mathbb{Q}_{n}}\left[\exp\left(\frac{\tilde{p}}{\tilde{p}-1}\int_{0}^{t}|f^{n}_{r}(X(s\wedge\tau_{n}))|\,\text{\rm{d}}s\right)\bigg{|}\mathscr{F}_{0}\right]\right)^{\frac{\tilde{p}-1}{\tilde{p}}}$
@@
By Lemma 4.3 and (4.16), we have the following corollary which is a key in the proofs of Theorem 2.1 and 2.2.
@@ 1. $(1)$
    $\mathbb{E}\left[\exp\left(\int_{s}^{t}|f_{u}(X(u))|\|X_{u}\|_{\tau}^{\alpha}\,\text{\rm{d}}u\right)\bigg{|}\mathscr{F}_{s}\right]\leq\exp\left[\varepsilon{\|X_{s}\|}_{\tau}^{2\alpha}+k\left(1+\|f\|^{\theta}_{\tilde{L}^{p}_{q}(s,t)}\right)\right],$ (4.24)
[…]
@@ 2. $(2)$
By (2.2), for any $n\geq 1$, $b^{(1)}_{n}$ is bounded and satisfies $(A_{3}^{\prime})$. So, Lemma 4.1 ensures that the SDE
@@
For $\lambda\geq\lambda_{0}$ and $u$ solving the PDE (4.7) such that (4.8) and (4.9) hold, let
@@
By $(A_{1})$$(A_{3})$, (4.9) and Itô’s formula, we find a constant $c_{1}>0$ such that
@@
Let $\Theta(t,x):=x+u(t,x)$ for $u$ solving the PDE (4.7) such that (4.8) and (4.9) holds, and denote
@@
By $(A_{1})$, $(A_{2})$, $(A_{3}^{\prime})$, (4.9) and Lemma 4.2, we find a constant $k_{0}>0$ such that
@@
According to Theorem 2.1 and Corollary 4.5, under the assumptions $(H_{1})$$(H_{2})$, for any $\mu\in\mathcal{C}_{k}^{\gamma}$, the SDE
@@ 5.1 Proof of (a)
It suffices to prove (5.3) for $k>0$ and large $N$. By Lemma 4.4, $(H_{1})$-$(H_{2})$ implies that for some constant $c_{1}>0$ we have
@@
$\sup_{\mu\in\mathcal{C}_{k}^{\gamma}}\mathbb{E}\bigg{(}\int_{0}^{T}|f_{0}(s,X^{\mu}(s))|^{2}\,\text{\rm{d}}s\bigg{)}^{k}<\infty.$
Combining this with $(H_{1})$-$(H_{2})$, $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{k,\text{\rm{e}}}^{\alpha}(\mathscr{C}_{\tau})$, the Itô formula and the BDG inequality, we find constants $c_{2},c_{3}>0$ such that
@@
Using Grönwall’s lemma, it holds
$\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right)\right]\leq c_{4}\text{\rm{e}}^{c_{4}T}\left(1+\int_{0}^{t}\text{\rm{e}}^{k\tau s}\|\mu_{s}\|_{k}^{k}\,\text{\rm{d}}s\right),\ \ t\in[0,T].$
Hence, by noting that $\mu\in\mathcal{C}^{\gamma,N}_{k}$, we can find a constant $c_{5}$ such that
$\begin{split}&\mathbb{E}\left[1+\|X_{t}^{\mu}\|_{\tau}^{k}\right]\leq 1+\text{\rm{e}}^{-k\tau t}\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right)\right]\\ &\leq c_{5}+c_{5}\int_{0}^{t}\text{\rm{e}}^{-k\tau(t-s)}\|\mu_{s}\|_{k}^{k}\,\text{\rm{d}}s\\ &\leq c_{5}+c_{5}N\text{\rm{e}}^{Nt}\int_{0}^{t}\text{\rm{e}}^{-N(t-s)}\text{\rm{d}}s\leq 2c_{5}\text{\rm{e}}^{Nt},\ \ t\in[0,T].\end{split}$ (5.6)
@@
$\begin{split}&\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\left(1+\|\Phi_{t}^{\gamma}\mu\|_{k}^{k}\right)=\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\mathbb{E}\left(1+\|X_{t}^{\mu}\|_{\tau}^{k}\right){\leq N},\quad N\geq N_{0},\ \mu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$
@@
$\begin{split}&U_{t}:=\mathbb{E}\left[\sup_{s\in[0,t]}\left(1+\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right)\right]\leq c_{6}+c_{6}\mathbb{E}\left(\int_{0}^{t}\left\{\text{\rm{e}}^{2\tau s}\|X_{s}^{\mu}\|_{\tau}^{2}+\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\right\}\,\text{\rm{d}}s\right)^{\frac{k}{2}},\\ &\leq c_{7}+c_{7}\left(\int_{0}^{t}\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\,\text{\rm{d}}s\right)^{\frac{k}{2}}+c_{7}\mathbb{E}\left\{\left[\sup_{s\in[0,t]}\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\right]^{1-\frac{k}{2}}\left(\int_{0}^{t}\text{\rm{e}}^{k\tau s}\|X_{s}^{\mu}\|_{\tau}^{k}\,\text{\rm{d}}s\right)^{\frac{k}{2}}\right\}\\ &\leq c_{7}+c_{7}\left(\int_{0}^{t}\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\,\text{\rm{d}}s\right)^{\frac{k}{2}}+\frac{1}{2}U_{t}+c_{8}\int_{0}^{t}U_{s}\,\text{\rm{d}}s,\ \ t\in[0,T],\end{split}$
then Grönwall’s lemma implies that
$U_{t}\leq 2c_{7}\text{\rm{e}}^{2c_{8}T}\left\{1+\left(\int_{0}^{t}\text{\rm{e}}^{2\tau s}\|\mu_{s}\|_{k}^{2}\,\text{\rm{d}}s\right)^{\frac{k}{2}}\right\},\ \ t\in[0,T].$
Thus, there exist constants $c_{9},c_{10}>0$ such that for any $\mu\in\mathcal{C}^{\gamma,N}_{k}$,
$\begin{split}&\mathbb{E}\left(1+\|X_{t}^{\mu}\|_{\tau}^{k}\right)\leq 1+\text{\rm{e}}^{-k\tau t}U_{t}\leq c_{9}+c_{9}\left(\int_{0}^{t}\text{\rm{e}}^{-2\tau(t-s)}\|\mu_{s}\|_{k}^{2}\,\text{\rm{d}}s\right)^{\frac{k}{2}}\\ &\leq c_{9}+c_{9}Ne^{Nt}\left(\int_{0}^{t}\text{\rm{e}}^{-2N(t-s)}\,\text{\rm{d}}s\right)^{\frac{k}{2}}\leq c_{9}+c_{10}N^{1-\frac{k}{2}}\text{\rm{e}}^{Nt},\ \ t\in[0,T].\end{split}$ (5.7)
@@
$\begin{split}&\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\left(1+\|\Phi_{t}^{\gamma}\mu\|_{k}^{k}\right)=\sup_{t\in[0,T]}\text{\rm{e}}^{-Nt}\mathbb{E}\left(1+\|X_{t}^{\mu}\|_{\tau}^{k}\right)\\ &\leq c_{9}+c_{10}N^{1-\frac{k}{2}}\leq N,\ \ \mu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$
@@ 5.2 Proof of (b)
By Theorem 2.1, $(H_{1})$-$(H_{2})$, (5.3) and $\gamma:=\mathscr{L}_{X_{0}}\in\mathscr{P}_{k}(\mathscr{C}_{\tau})$, we find a constant $C_{0}(N)>0$ such that
@@
Next, since $\mu,\nu\in\mathcal{C}^{\gamma,N}_{k}$, we find a constant $C_{0}(N)>0$ such that
$\sup_{t\in[0,T]}\mathbb{W}_{k,var}(\mu_{t},\nu_{t})\leq C_{0}(N).$
Thus, by $(H_{1})$ and $(H_{2})$, we find a constant $C_{1}(N)>0$ such that for any $\mu,\nu\in\mathcal{C}^{\gamma,N}_{k}$,
@@
Recalling that $H\in L_{loc}^{1}([0,\infty);(0,\infty))$, we have
$\int_{0}^{t}H(s)\left(1\wedge\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\right)\,\text{\rm{d}}s\leq\int_{0}^{t}H(s)\,\text{\rm{d}}s<\infty,$
then by $(H_{2})$ and Girsanov’s theorem, for any $t\in[0,T]$,
@@
$\begin{split}&\mathbb{E}\left(|R_{t}-1|^{2}\big{|}\mathscr{F}_{0}\right)=\mathbb{E}\left[\text{\rm{e}}^{2\int_{0}^{t}\langle\zeta_{s},\text{\rm{d}}W(s)\rangle-2\int_{0}^{t}|\zeta_{s}|^{2}\,\text{\rm{d}}s+\int_{0}^{t}|\zeta_{s}|^{2}\,\text{\rm{d}}s}\bigg{|}\mathscr{F}_{0}\right]-1\\ &\leq\text{\rm{e}}^{C_{1}(N)\int_{0}^{t}H(s)\left(1\wedge\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\right)\,\text{\rm{d}}s}-1\leq C_{2}(N)\int_{0}^{t}H(s)\mathbb{W}_{k,var}(\mu_{s},\nu_{s})^{2}\,\text{\rm{d}}s\\ &\leq C_{3}(N)\text{\rm{e}}^{2\theta t}\tilde{\mathbb{W}}_{k,\theta,var}(\mu,\nu)^{2}\int_{0}^{t}H(s)\text{\rm{e}}^{-2\theta(t-s)}\,\text{\rm{d}}s,\ \ \mu,\nu\in\mathcal{C}^{\gamma,N}_{k}.\end{split}$ (5.11)
@@
where $\mathscr{L}_{X_{t}^{\mu}|\mathbb{Q}_{t}}$ stands for the distribution of $X_{t}^{\mu}$ under $\mathbb{Q}_{t}$. Thus, by (5.8), (5.9), (LABEL:N1) and Hölder’s inequality, we find constant $C_{4}(N)>0$ such that
@@
Noting that $H\in L_{loc}^{1}([0,\infty);(0,\infty))$ yields that
@@
Let $k>1$. Let $\Theta(t,\cdot):=id+u(t,\cdot)$ for $u$ solving (4.7) such that (4.8) and (4.9) holds. Let
@@
$\text{\rm{d}}X^{\nu}(t)=b(t,X_{t}^{\nu},\nu_{t})\text{\rm{d}}t+\sigma(t,X^{\nu}_{t})\text{\rm{d}}W(t),\ \ t\in[0,T]$ (6.3)
@@
Combining this with (4.12), Itô’s isometry and Young’s inequality, we find a constant $c_{1}>0$ such that
@@
$\bar{\zeta}_{s}:=\left\{\sigma^{*}(\sigma\sigma^{*})^{-1}\right\}(s,X^{\mu}_{s})\left[b(s,X^{\mu}_{s},\mu_{s})-b(s,X^{\mu}_{s},\nu_{s})\right],$
@@
$|\bar{\zeta}_{s}|^{2}\leq c_{1}K_{2}\mathbb{W}_{2}(\mu_{s},\nu_{s})^{2}\leq c_{2}K_{2}\text{\rm{e}}^{c_{2}s}\mathbb{W}_{2}(\mu,\nu)^{2}$ (6.6)
@@
$\text{\rm{d}}X^{\mu}(s)=b(s,X^{\mu}_{s},\nu_{s})\text{\rm{d}}s+\sigma(s,X^{\mu}_{s})\text{\rm{d}}\bar{W}(s),\ \ \mathscr{L}_{X^{\mu}_{0}}=\mu,\ s\in[0,t].$ (6.7)
@@
$\text{\rm{d}}Y(s)=$$\{b(s,Y_{s},\nu_{s})+\kappa\sigma(s,Y_{s})\sigma(s,X^{\mu}_{s})^{-1}(X^{\mu}(s)-Y(s))\}\text{\rm{d}}s$ (6.8)
$+\sigma(s,Y_{s})\text{\rm{d}}\bar{W}(s),\ \ s\in[0,t],\ \ Y_{0}=X_{0}^{\nu}.$
@@
$\tilde{\zeta}_{s}:=\kappa\sigma(s,X^{\mu}_{s})^{-1}(X^{\mu}(s)-Y(s)),$
@@
$\text{\rm{d}}Y(s)=b(s,Y_{s},\nu_{s})\text{\rm{d}}s+\sigma(s,Y_{s})\text{\rm{d}}\tilde{W}(s),\ \ s\in[0,t],\ \ Y_{0}=X_{0}^{\nu},$ (6.9)
which together with the uniqueness of (6.3) derives that $\mathscr{L}_{Y_{t}|\mathbb{Q}}=\mathscr{L}_{X_{t}^{\nu}}$. Moreover, if we choose $\mathscr{F}_{0}$-measurable random variables $X_{0}^{\mu}$ and $X_{0}^{\nu}$ on $\mathscr{C}_{\tau}$ such that
@@
By (6.7), $(H^{\prime})$ and the proof of [5, Lemma 3.3], for any $p>0$ and $\tau_{0}\in(0,\tau)$ we find a constant $\kappa>\tau$ to define $Y_{s}$ in (6.8) such that
@@
holds for some constant $c>0$. Therefore, applying Young’s inequality in [2, Lemma 2.4],
@@
$\leq c_{3}K_{2}\text{\rm{e}}^{c_{3}t}\mathbb{W}_{2}(\mu,\nu)^{2}+c_{3}\int_{0}^{t}\mathbb{E}_{\mathbb{Q}}\|X^{\mu}_{s}-Y_{s}\|_{\tau}^{2}\,\text{\rm{d}}s\leq\left(c_{3}K_{2}\text{\rm{e}}^{c_{3}t}+c_{4}\right)\mathbb{W}_{2}(\mu,\nu)^{2}.$
@@
Acknowledgements
The authors would like to thank the associated editor and referees for their helpful comments and suggestions.